# Pay Less Attention to Function Words for Free Robustness of Vision-Language Models

### v2

@@ Appendix C Full results for untargeted attacks
Appendix D Full results for ablation studies
@@
Appendix E Details for function word dictionary
We provide the function word dictionary we used as follows: “am, is, are, was, were, be, been, being, have, has, had, do, does, did, will, would, shall, should, may, might, must, can, could, ought, dare, need, used, to, a, an, the, and, but, if, or, because, as, until, while, of, at, by, for, with, about, against, between, into, through, during, before, after, above, below, to, from, in, out, on, off, over, under, again, further, then, once, here, there, when, where, why, how, all, any, both, each, few, more, most, other, some, such, no, nor, not, only, own, same, so, than, too, very”.
Appendix F Visualization of attention scores

### v3

@@ Appendix D Full results for untargeted attacks
Appendix E Full results for ablation studies
@@
Appendix F Details for function word dictionary
We provide the function word dictionary we used as follows: “am, is, are, was, were, be, been, being, have, has, had, do, does, did, will, would, shall, should, may, might, must, can, could, ought, dare, need, used, to, a, an, the, and, but, if, or, because, as, until, while, of, at, by, for, with, about, against, between, into, through, during, before, after, above, below, from, in, out, on, off, over, under, again, further, then, once, here, there, when, where, why, how, all, any, both, each, few, more, most, other, some, such, no, nor, not, only, own, same, so, than, too, very”.
Appendix G Visualization of attention scores